OpenStudy (anonymous):

. Identify a, b, and c for use in the Quadratic Formula. 5x^2 =4x - 7 A. a = 5, b = 4, c = -7 B. a = 5, b = -4, c = 7 C. a = -5, b = -4, c = -7 D. a = 4, b = -7, c = 5

hartnn (hartnn):

bring all terms on one side to get the form \(ax^2+bx+c=0\) so, \(5x^2-4x+7=0\) compare it with \(ax^2+bx+c=0\)

hartnn (hartnn):

 now can you find a,b,c ?